2x-y+z=1
x+2y+3z=8
3x+y-4z=1

Answer by Edwin McCravy(20062)

2x- y+ z=1
 x+2y+3z=8
3x+ y-4z=1

There are many ways to solve this.  
Since you didn't state a method, we're
not just going to guess one particular
way.  If you want to see it solved, 
specify which method you are studying,
tell me in the thank-you note form below,
and I'll get back to you.

The solution is

x=1, y=2, z=1

Edwin